OpenStudy (anonymous):

what is the inflection point of the equation x^4-6x^2+3? (a)3 (b)2 (c)1 (d)0 (e)√3

OpenStudy (anonymous):

Diff. the equations two times and put it equal to zero. Can you do it?

OpenStudy (alexwee123):

just take the double derivative and set it to 0

OpenStudy (alexwee123):

f(x)= x^4-6x^2+3 f'(x)=4x^3 -12x f''(x)=12x^2-12=0 12x^2=12 x^2=1 x=1 so your answer is C

OpenStudy (anonymous):

thankss you guys :)